# Bell Instability and Cosmic-Ray Acceleration in AGN Ultrafast Outflow Shocks

### v2

@@ Abstract
1 INTRODUCTION
Cosmic rays (CRs) are high-energy charged particles propagating through interstellar and intergalactic space, spanning energies from $10^{9}\,\mathrm{eV}$ to $10^{20}\,\mathrm{eV}$. Understanding their origin is not only a fundamental problem in high-energy astrophysics but also essential for galaxy evolution. Despite extensive studies, the origin of CRs across different energy ranges remains unresolved. Supernova remnants (SNRs) are regarded as the most promising candidates up to the knee energy (a few $\times 10^{15}\,\mathrm{eV}$) (1934PhRv...46...76B; 1978MNRAS.182..147B; 1978ApJ...221L..29B), whereas the sources and acceleration mechanisms at and above the knee energy are largely unknown.
Active galactic nuclei (AGN) are among the leading candidates for the origin of high-energy CRs across and beyond the knee energy. Proposed acceleration sites in AGN include relativistic jets (2004NewAR..48..381A; 2014PhRvD..90b3007M; 2015SSRv..191..519S; 2018ApJ...863L..10A; 2018ApJ...865..124M), coronae (2020ApJ...891L..33I; 2020PhRvL.125a1101M; 2021ApJ...922...45K; 2022ApJ...939...43E; 2022ApJ...941L..17M), accretion flows (2019PhRvD.100h3014K; 2021A&A...649A..87G), and outflows (2016A&A...596A..68L; 2017PhRvD..95f3007W; 2018ApJ...858....9L; 2022arXiv220702097I; 2023MNRAS.526..181P). In addition, AGN are considered promising multi-messenger sources because the acceleration of CRs is expected to produce high-energy neutrinos and gamma rays (2017PhRvD..95f3007W; 2018ApJ...858....9L; 2020PhRvL.124e1103A; 2021ApJ...921..144A; 2022Sci...378..538I; 2022arXiv220702097I; 2025JCAP...07..013P; 2025ApJ...980..131S).
@@ 1 INTRODUCTION
UFOs colliding with ambient gas are expected to form shock structures. These shocks have been proposed as promising sites for diffusive shock acceleration (DSA) (1978MNRAS.182..147B; 1978ApJ...221L..29B; 1983RPPh...46..973D; 1987PhR...154....1B; 2016RPPh...79d6901M), in which CRs repeatedly cross the shock front via pitch-angle scattering off magnetic turbulence, gaining energy efficiently. Recent studies indicate that protons may be accelerated up to $10^{18}\,\mathrm{eV}$ and heavy nuclei up to $10^{20}\,\mathrm{eV}$ at reverse shocks of the UFOs (2023MNRAS.526..181P; 2025MNRAS.539.2435E).
@@
The nonresonant hybrid instability (NRH instability, also known as the Bell instability) offers a self-consistent mechanism for determining key magnetic-field parameters such as $\varepsilon_{B}$ and $\delta B/B_{0}$ in the shock vicinity (2004MNRAS.353..550B). This instability is driven by the CR current escaping upstream of the shock, which induces a return current in the background plasma and amplifies circularly polarized Alfvén waves. Magnetic-field amplification regions have been observed in SNRs (2003ApJ...584..758V; 2003ApJ...589..827B; 2005ApJ...621..793B), and these are often interpreted as possible evidence of NRH instability.
@@
NRH instability has been investigated both analytically and numerically. Analytically, 2004MNRAS.353..550B derived the linear growth theory. Numerically, a wide range of methods has been applied, including full particle-in-cell (PIC) simulations (2015PhRvL.114h5003P), kinetic hybrid approaches (2014ApJ...783...91C; 2014ApJ...794...46C), MHD-PIC methods (2012ApJ...759...73N), and hybrid MHD–Vlasov-Fokker-Planck (VFP) models (2013MNRAS.431..415B; 2013MNRAS.430.2873R). While PIC simulations capture kinetic effects from shock formation to nonthermal injection, they are computationally prohibitive for tracing particle acceleration orders of magnitude larger than the injection. Hybrid MHD–VFP models reduce computational costs by expanding the CR Boltzmann equation into multipoles, but these studies in most cases fixed the CR current, preventing a fully self-consistent treatment of particle acceleration and field amplification.
@@
2019ApJ...872...46I; 2021ApJ...922....7I; 2024ApJ...965..113I developed a numerical code capable of simultaneously evolving the background plasma and CR current in one-dimensional setups. Their results demonstrated that finite spatial extent of the upstream CR current supresses the growth of NRH instability compared to previous theoretical models that assume infinite spatial extent of the CR current. Nevertheless, under favorable conditions, NRH instability can amplify magnetic fields sufficiently to enable $10^{15}~\mathrm{eV}$ acceleration in very early stage of young SNRs.
@@
In this work, we employ the numerical code of 2019ApJ...872...46I to investigate the growth and saturation of NRH instability in the context of UFOs. By systematically varying control parameters such as background magnetic field $B_{0}$, injection rate $\eta$, and initial amplitude of the magnetic fluctuation $\xi_{B,\mathrm{ini}}$ defined by Eq. (26), we assess whether the maximum CR energy $E_{\max}$ in UFO shocks can be determined self-consistently, without resorting to phenomenological parameters.
@@
The structure of this paper is as follows. In Sec. 2, we describe the governing MHD–CR Boltzmann equations, diffusion coefficient, and the adopted boundary and initial conditions, including $B_{0}$, $\xi_{B,\mathrm{ini}}$, $\eta$, and $\mathrm{p}\gamma$ cooling. In Sec. 3, we present the main results: (i) simulations without the NRH term reproduce test-particle DSA, (ii) in weak-field regimes ($B_{0}=10^{-5}$$10^{-4}\,\mathrm{G}$), NRH instability amplifies the magnetic field, driving $E_{\max}$ and $\varepsilon_{B}$ to converge regardless of $\xi_{B,\mathrm{ini}}$, (iii) in strong-field regimes ($B_{0}\gtrsim 10^{-3}\,\mathrm{G}$), NRH instability is ineffective and $E_{\max}$ depends on the initial conditions, (iv) variations in interstellar medium (ISM) density affect acceleration efficiency, and (v) $\mathrm{p}\gamma$ cooling can limit the maximum energy.
2 Basic Equations and Simulation Setup
@@
To investigate particle acceleration in UFOs, we adopt the following physical assumptions. More detailed initial conditions are presented in Sec. 2.2.
@@
    When a UFO collides with the ISM, it generates a shock structure composed of a reverse shock, contact discontinuity, and forward shock. In this study, we focus on the local region around the reverse shock and analyze particle acceleration there (see Fig. 2). This choice is motivated by previous studies that proposed the UFO reverse shock as a promising site for acceleration up to EeV energies (2023MNRAS.526..181P; 2025MNRAS.539.2435E). Moreover, in the upstream rest frame, the forward shock decelerates with time, whereas the reverse shock remains comparable to the wind velocity, so we may expect more favorable for efficient particle acceleration.
@@
    The wind is assumed to be steady and uniform [^1] Variability of UFOs has been observed on timescales from months down to days (2008MNRAS.385L.108R; 2009A&A...504..401C; 2009MNRAS.397..249P). Moreover, both theoretical and observational studies suggest that UFOs are clumpy and inhomogeneous (2013PASJ...65...88T; 2025Natur.641.1132X). These factors may affect particle acceleration and will be considered in future work..
@@
    and parallel to the shock normal over a coherence length $\ell_{\mathrm{coh}}$. In addition, broadband, circular-polarized Alfvén wave turbulence is superposed on the $y$ and $z$ components of magnetic field (see Sec. 2.2.5 for detail). Although Parker-type wind solutions imply that the large scale magnetic field becomes toroidally dominated at large radii, UFO outflows are expected to be intermittent and turbulent rather than laminar. In such a turbulent dynamo, the magnetic field is amplified on the hydrodynamic eddy scale and naturally forms local patches with a coherent mean field. The NRH instability can grow when the coherence length of $B_{x}$ exceeds the wavelength of the NRH mode with maximum growth, as given by Eq. (28), that is, when $\ell_{\mathrm{coh}}>\lambda_{\mathrm{NRH}}$. Even if the large scale geometry of the UFO reverse shock is quasi-perpendicular, such locally quasi-parallel patches with a coherent background magnetic field are expected to exist in a turbulent dynamo. In our setup, the initial condition $\delta B/B_{0}<1$ is imposed at the NRH maximum growth scale, so that the NRH mode effectively samples a nearly uniform background magnetic field with $\ell_{\mathrm{coh}}>\lambda_{\mathrm{NRH}}$, while the field can still be strongly tangled on larger hydrodynamic scales.
@@
2.1 Basic Equations
In this study, we solve the coupled system of magnetohydrodynamic (MHD) equations for the fluid component and telegraph-type diffusion–convection equations for the CR distribution function (2013MNRAS.431..415B; 2019ApJ...872...46I; 2021ApJ...922....7I; 2024ApJ...965..113I). The MHD equations are expressed as follows. The continuity equation is given by
@@ 2.1 Basic Equations
[…]
@@
[…]
@@
Here, $Q_{\mathrm{inj}}$ is the CR injection rate, $t_{\mathrm{p}\gamma}(p)$ is the timescale of $\mathrm{p}\gamma$ interactions near the AGN source (see Fig. 1), and $D_{\|}(p,B)$ is the diffusion coefficient. In this work, we incorporate the $\mathrm{p}\gamma$ energy-loss timescale derived by 2023MNRAS.526..181P into the numerical framework of 2019ApJ...872...46I. Note that Eqs. (13) and (14) are reduced to the diffusion convection equation in the limit $c\rightarrow\infty$ (see, 2021ApJ...922....7I for numerical tests).
The injection rate $Q_{\mathrm{inj}}$ is modeled as follows. A fraction $\eta$ of upstream fluid particles with momentum $p_{\text{inj}}$ are assumed to be injected into the acceleration process (10.1111/j.1365-2966.2005.09227.x):
@@
with $m_{\mathrm{g}}$ the mean gas mass and $T_{\text{sh}}$ the downstream temperature. For strong shocks with $\gamma=5/3$, the Rankine–Hugoniot relation gives $T_{\text{sh}}=3m_{\text{g}}v_{\text{sh}}/(16k_{\text{B}})$. We adopt $m_{\text{g}}=1.27m_{\text{p}}$, consistent with the solar composition inferred from Cygnus A emission-line ratios (2006agna.book.....O, Tab. 13.4). For numerical reasons, the momentum range of CR is restricted to $p_{\min}>p_{\mathrm{inj}}$ up to $p_{\max}$. Between $p_{\mathrm{inj}}$ and $p_{\min}$, we assume the standard DSA spectrum $f_{0}(x_{\text{sh}})\propto p^{-4}$ (2019ApJ...872...46I; 2021ApJ...922....7I; 2024ApJ...965..113I). Accordingly, the injection rate used in simulations is rewritten as
@@
Figure 1: Comparison of $\mathrm{p}\gamma$ cooling and acceleration timescales as a function of proton energy. The blue solid line shows the $\mathrm{p}\gamma$ cooling timescale, the orange dashed line indicates the simulation runtime (the propagation timescale of the reverse shock across pc scales), and the red dot–dashed line represents the acceleration timescale described in Sec. 3.3, calculated from Eq. (36) with $\xi_{B}=0.1$, $B_{0}=10^{-2}~\mathrm{G}$, and $v_{\text{sh}}=5.0\times 10^{9}~\mathrm{cm~s^{-1}}$ (see Tab. 1 for the fiducial model).
@@
The validity of Eq. (18) is supported by two arguments. First, kinetic hybrid simulations (2014ApJ...794...47C) and test-particle calculations (2016APh....73....1R) have shown that NRH-amplified fields drive diffusion close to the Bohm-like limit. Second, NRH instability alone amplifies only small-scale magnetic fluctuations, with wavelengths much shorter than the gyroradius at $E_{\max}$. Filamentation instability can transfer this energy to larger scales, amplifying turbulence up to the gyroradius scale (2012MNRAS.419.2433R; 2013ApJ...765L..20C).
2.2 Initial Conditions
@@
2.2.1 Global Fluid Initial Conditions
The wind launched from a supermassive black hole, including UFOs, is expected to form a bubble structure similar to stellar winds, as suggested by both theory and observations (see Fig. 2 left) (2012ASPC..460..235Z; 2018MNRAS.474.3673R; 2018ApJ...867...88R; 2018ApJ...858....9L; 2019A&A...628A.118B; 2023ApJ...951..146M). We performed one-dimensional MHD simulations of a Riemann problem, where a stationary homogeneous ISM is placed on the left side of $x=0$ and a leftward-propagating wind is injected on the right side. The time evolution of this system generates a shock structure consisting of a reverse shock, contact discontinuity, and forward shock. The resulting density distribution of the bubble is shown in Fig. 2 (left). The detailed quantative conditions for the left and right states of the Riemann problem are given in Sec. 2.2.3 and 2.2.4. Solving Eqs. (13) and (14) for the whole system is computationally expensive. To reduce it, we extract only the local region around the reverse shock and solve Eqs. (13) and (14) in couple with the Bell-MHD Eqs. (2)-(9) (see, right panel of Fig. 2).
2.2.2 Region Around the Reverse Shock
The simulation domain is set to $L_{\text{box}}=10~\text{pc}$ with a runtime of $t_{\text{final}}=100~\text{yr}$. [^2] For the fiducial model with $B_{0}=10^{-5}~\mathrm{G}$, the box size is increased to $L_{\text{box}}=15~\text{pc}$ to prevent the anisotropic CR current (Eq. (11)) from reaching the boundary. Typical wind velocities are $v_{\text{wind}}\sim 0.1c$$0.4c$ (2003MNRAS.345..705P; 2005A&A...442..461D; 2006ApJ...646..783M; 2006AN....327.1012C; 2007ApJ...670..978B; 2009A&A...504..401C; 2009ApJ...701..493R; 2010A&A...521A..57T; 2015MNRAS.451.4169G; 2015Sci...347..860N; 2019MNRAS.482.5316M; 2021NatAs...5...13L; 2025Natur.641.1132X). This setup corresponds to the propagation timescale of a UFO traveling about 1 pc from the central black hole. Observationally, UFOs are typically detected at distances $10^{18}~\text{cm}$ (2012MNRAS.422L...1T; 2015ARA&A..53..115K; 2019GReGr..51...65Z), placing the 1 pc scale within the observed distribution range.
The wind and ISM parameters are summarized in Tab. 1. The mass outflow rate is set to $\dot{M}\sim 0.01$$1~M_{\odot}/\mathrm{yr}$ (2012MNRAS.422L...1T; 2015MNRAS.451.4169G; 2019ApJ...871..156M; 2021NatAs...5...13L).
2.2.3 Wind Region

### v3

# Bell Instability and Cosmic-Ray Acceleration in Active Galactic Nuclei Ultrafast Outflow Shocks

@@ Abstract
I INTRODUCTION
Cosmic rays (CRs) are high-energy charged particles propagating through interstellar and intergalactic space, spanning energies from $10^{9}\,\mathrm{eV}$ to $10^{20}\,\mathrm{eV}$. Understanding their origin is not only a fundamental problem in high-energy astrophysics but also essential for galaxy evolution. Despite extensive studies, the origin of CRs across different energy ranges remains unresolved. Supernova remnants (SNRs) are regarded as the most promising candidates up to the knee energy (a few $\times 10^{15}\,\mathrm{eV}$) (Baade and Zwicky, 1934; Bell, 1978; Blandford and Ostriker, 1978), whereas the sources and acceleration mechanisms at and above the knee energy are largely unknown.
Active galactic nuclei (AGN) are among the leading candidates for the origin of high-energy CRs across and beyond the knee energy. Proposed acceleration sites in AGN include relativistic jets (Atoyan and Dermer, 2004; Murase et al., 2014; Sironi et al., 2015; Ansoldi et al., 2018; Murase et al., 2018), coronae (Inoue et al., 2020; Murase et al., 2020; Kheirandish et al., 2021; Eichmann et al., 2022; Murase, 2022), accretion flows (Kimura et al., 2019; Gutiérrez et al., 2021), and outflows (Lamastra et al., 2016; Wang and Loeb, 2017; Liu et al., 2018; Inoue et al., 2022; Peretti et al., 2023). In addition, AGN are considered promising multi-messenger sources because the acceleration of CRs is expected to produce high-energy neutrinos and gamma rays (Wang and Loeb, 2017; Liu et al., 2018; Aartsen et al., 2020; Ajello et al., 2021; IceCube Collaboration et al., 2022; Inoue et al., 2022; Peretti et al., 2025; Sakai et al., 2025).
@@ I INTRODUCTION
UFOs colliding with ambient gas are expected to form shock structures. These shocks have been proposed as promising sites for diffusive shock acceleration (DSA) (Bell, 1978; Blandford and Ostriker, 1978; Drury, 1983; Blandford and Eichler, 1987; Marcowith et al., 2016), in which CRs repeatedly cross the shock front via pitch-angle scattering off magnetic turbulence, gaining energy efficiently. Recent studies indicate that protons may be accelerated up to $10^{18}\,\mathrm{eV}$ and heavy nuclei up to $10^{20}\,\mathrm{eV}$ at reverse shocks of the UFOs (Peretti et al., 2023; Ehlert et al., 2025).
@@
The nonresonant hybrid instability (NRH instability, also known as the Bell instability) offers a self-consistent mechanism for determining key magnetic-field parameters such as $\varepsilon_{B}$ and $\delta B/B_{0}$ in the shock vicinity (Bell, 2004). This instability is driven by the CR current escaping upstream of the shock, which induces a return current in the background plasma and amplifies circularly polarized Alfvén waves. Magnetic-field amplification regions have been observed in SNRs (Vink and Laming, 2003; Bamba et al., 2003, 2005), and these are often interpreted as possible evidence of NRH instability.
@@
NRH instability has been investigated both analytically and numerically. Analytically, Bell (2004) derived the linear growth theory. Numerically, a wide range of methods has been applied, including full particle-in-cell (PIC) simulations (Park et al., 2015), kinetic hybrid approaches (Caprioli and Spitkovsky, 2014a, b), MHD-PIC methods (Niemiec et al., 2012), and hybrid MHD–Vlasov-Fokker-Planck (VFP) models (Bell et al., 2013; Reville and Bell, 2013). While PIC simulations capture kinetic effects from shock formation to nonthermal injection, they are computationally prohibitive for tracing particle acceleration orders of magnitude larger than the injection. Hybrid MHD–VFP models reduce computational costs by expanding the CR Boltzmann equation into multipoles, but these studies in most cases fixed the CR current, preventing a fully self-consistent treatment of particle acceleration and field amplification.
@@
Inoue (2019); Inoue et al. (2021, 2024) developed a numerical code capable of simultaneously evolving the background plasma and CR current in one-dimensional setups. Their results demonstrated that finite spatial extent of the upstream CR current supresses the growth of NRH instability compared to previous theoretical models that assume infinite spatial extent of the CR current. Nevertheless, under favorable conditions, NRH instability can amplify magnetic fields sufficiently to enable $10^{15}~\mathrm{eV}$ acceleration in very early stage of young SNRs.
@@
In this work, we employ the numerical code of Inoue (2019) to investigate the growth and saturation of NRH instability in the context of UFOs. By systematically varying control parameters such as background magnetic field $B_{0}$, injection rate $\eta$, and initial amplitude of the magnetic fluctuation $\xi_{B,\mathrm{ini}}$ defined by Eq. (26), we assess whether the maximum CR energy $E_{\max}$ in UFO shocks can be determined self-consistently, without resorting to phenomenological parameters.
@@
The structure of this paper is as follows. In Sec. II, we describe the governing MHD–CR Boltzmann equations, diffusion coefficient, and the adopted boundary and initial conditions, including $B_{0}$, $\xi_{B,\mathrm{ini}}$, $\eta$, and $\mathrm{p}\gamma$ cooling. In Sec. III, we present the main results: (i) simulations without the NRH term reproduce test-particle DSA, (ii) in weak-field regimes ($B_{0}=10^{-5}$$10^{-4}\,\mathrm{G}$), NRH instability amplifies the magnetic field, driving $E_{\max}$ and $\varepsilon_{B}$ to converge regardless of $\xi_{B,\mathrm{ini}}$, (iii) in strong-field regimes ($B_{0}\gtrsim 10^{-3}\,\mathrm{G}$), NRH instability is ineffective and $E_{\max}$ depends on the initial conditions, (iv) variations in interstellar medium (ISM) density affect acceleration efficiency, and (v) $\mathrm{p}\gamma$ cooling can limit the maximum energy.
II Basic Equations and Simulation Setup
@@
To investigate particle acceleration in UFOs, we adopt the following physical assumptions. More detailed initial conditions are presented in Sec. II.2.
@@
    When a UFO collides with the ISM, it generates a shock structure composed of a reverse shock, contact discontinuity, and forward shock. In this study, we focus on the local region around the reverse shock and analyze particle acceleration there (see Fig. 2). This choice is motivated by previous studies that proposed the UFO reverse shock as a promising site for acceleration up to EeV energies (Peretti et al., 2023; Ehlert et al., 2025). Moreover, in the upstream rest frame, the forward shock decelerates with time, whereas the reverse shock remains comparable to the wind velocity, so we may expect more favorable for efficient particle acceleration.
@@
    The wind is assumed to be steady and uniform [^1] Variability of UFOs has been observed on timescales from months down to days (Reeves et al., 2008; Cappi et al., 2009; Pounds and Reeves, 2009). Moreover, both theoretical and observational studies suggest that UFOs are clumpy and inhomogeneous (Takeuchi et al., 2013; Xrism Collaboration et al., 2025). These factors may affect particle acceleration and will be considered in future work..
@@
    and parallel to the shock normal over a coherence length $\ell_{\mathrm{coh}}$. In addition, broadband, circular-polarized Alfvén wave turbulence is superposed on the $y$ and $z$ components of magnetic field (see Sec. II.2.5 for detail). Although Parker-type wind solutions imply that the large scale magnetic field becomes toroidally dominated at large radii, UFO outflows are expected to be intermittent and turbulent rather than laminar. In such a turbulent dynamo, the magnetic field is amplified on the hydrodynamic eddy scale and naturally forms local patches with a coherent mean field. The NRH instability can grow when the coherence length of $B_{x}$ exceeds the wavelength of the NRH mode with maximum growth, as given by Eq. (28), that is, when $\ell_{\mathrm{coh}}>\lambda_{\mathrm{NRH}}$. Even if the large scale geometry of the UFO reverse shock is quasi-perpendicular, such locally quasi-parallel patches with a coherent background magnetic field are expected to exist in a turbulent dynamo. In our setup, the initial condition $\delta B/B_{0}<1$ is imposed at the NRH maximum growth scale, so that the NRH mode effectively samples a nearly uniform background magnetic field with $\ell_{\mathrm{coh}}>\lambda_{\mathrm{NRH}}$, while the field can still be strongly tangled on larger hydrodynamic scales.
@@
II.1 Basic Equations
In this study, we solve the coupled system of magnetohydrodynamic (MHD) equations for the fluid component and telegraph-type diffusion–convection equations for the CR distribution function (Bell et al., 2013; Inoue, 2019; Inoue et al., 2021, 2024). The MHD equations are expressed as follows. The continuity equation is given by
@@ II.1 Basic Equations
$\frac{\partial\epsilon}{\partial t}+\frac{\partial}{\partial x}\left\{v_{x}\left(\epsilon+P+\frac{B_{y}^{2}+B_{z}^{2}}{8\pi}\right)-B_{x}\frac{\bm{B}\cdot\bm{v}}{4\pi}\right\}=0.$ (6)
@@
The CR distribution function is expanded with respect to the pitch-angle cosine relative to the background magnetic field $\bm{B}_{0}$ as follows,
@@
Here, $Q_{\mathrm{inj}}$ is the CR injection rate, $t_{\mathrm{p}\gamma}(p)$ is the timescale of $\mathrm{p}\gamma$ interactions near the AGN source (see Fig. 1), and $D_{\|}(p,B)$ is the diffusion coefficient. In this work, we incorporate the $\mathrm{p}\gamma$ energy-loss timescale derived by Peretti et al. (2023) into the numerical framework of Inoue (2019). Note that Eqs. (13) and (14) are reduced to the diffusion convection equation in the limit $c\rightarrow\infty$ (see, Inoue et al. (2021) for numerical tests).
The injection rate $Q_{\mathrm{inj}}$ is modeled as follows. A fraction $\eta$ of upstream fluid particles with momentum $p_{\text{inj}}$ are assumed to be injected into the acceleration process (Blasi et al., 2005a):
@@
with $m_{\mathrm{g}}$ the mean gas mass and $T_{\text{sh}}$ the downstream temperature. For strong shocks with $\gamma=5/3$, the Rankine–Hugoniot relation gives $T_{\text{sh}}=3m_{\text{g}}v_{\text{sh}}/(16k_{\text{B}})$. We adopt $m_{\text{g}}=1.27m_{\text{p}}$, consistent with the solar composition inferred from Cygnus A emission-line ratios (Osterbrock and Ferland, 2006, Tab. 13.4). For numerical reasons, the momentum range of CR is restricted to $p_{\min}>p_{\mathrm{inj}}$ up to $p_{\max}$. Between $p_{\mathrm{inj}}$ and $p_{\min}$, we assume the standard DSA spectrum $f_{0}(x_{\text{sh}})\propto p^{-4}$ (Inoue, 2019; Inoue et al., 2021, 2024). Accordingly, the injection rate used in simulations is rewritten as
@@
Figure 1: Comparison of $\mathrm{p}\gamma$ cooling and acceleration timescales as a function of proton energy. The blue solid line shows the $\mathrm{p}\gamma$ cooling timescale, the orange dashed line indicates the simulation runtime (the propagation timescale of the reverse shock across pc scales), and the red dot–dashed line represents the acceleration timescale described in Sec. III.3, calculated from Eq. (36) with $\xi_{B}=0.1$, $B_{0}=10^{-2}~\mathrm{G}$, and $v_{\text{sh}}=5.0\times 10^{9}~\mathrm{cm~s^{-1}}$ (see Tab. 1 for the fiducial model).
@@
The validity of Eq. (18) is supported by two arguments. First, kinetic hybrid simulations (Caprioli and Spitkovsky, 2014c) and test-particle calculations (Roh et al., 2016) have shown that NRH-amplified fields drive diffusion close to the Bohm-like limit. Second, NRH instability alone amplifies only small-scale magnetic fluctuations, with wavelengths much shorter than the gyroradius at $E_{\max}$. Filamentation instability can transfer this energy to larger scales, amplifying turbulence up to the gyroradius scale (Reville and Bell, 2012; Caprioli and Spitkovsky, 2013).
II.2 Initial Conditions
@@
II.2.1 Global Fluid Initial Conditions
The wind launched from a supermassive black hole, including UFOs, is expected to form a bubble structure similar to stellar winds, as suggested by both theory and observations (see Fig. 2 left) (Zubovas and King, 2012; Richings and Faucher-Giguère, 2018; Revalski et al., 2018; Liu et al., 2018; Bischetti et al., 2019; Maksym et al., 2023). We performed one-dimensional MHD simulations of a Riemann problem, where a stationary homogeneous ISM is placed on the left side of $x=0$ and a leftward-propagating wind is injected on the right side. The time evolution of this system generates a shock structure consisting of a reverse shock, contact discontinuity, and forward shock. The resulting density distribution of the bubble is shown in Fig. 2 (left). The detailed quantative conditions for the left and right states of the Riemann problem are given in Sec. II.2.3 and II.2.4. Solving Eqs. (13) and (14) for the whole system is computationally expensive. To reduce it, we extract only the local region around the reverse shock and solve Eqs. (13) and (14) in couple with the Bell-MHD Eqs. (2)-(9) (see, right panel of Fig. 2).
II.2.2 Region Around the Reverse Shock
The simulation domain is set to $L_{\text{box}}=10~\text{pc}$ with a runtime of $t_{\text{final}}=100~\text{yr}$. [^2] For the fiducial model with $B_{0}=10^{-5}~\mathrm{G}$, the box size is increased to $L_{\text{box}}=15~\text{pc}$ to prevent the anisotropic CR current (Eq. (11)) from reaching the boundary. Typical wind velocities are $v_{\text{wind}}\sim 0.1c$$0.4c$ (Pounds et al., 2003; Dadina et al., 2005; Markowitz et al., 2006; Cappi, 2006; Braito et al., 2007; Cappi et al., 2009; Reeves et al., 2009; Tombesi et al., 2010; Gofford et al., 2015; Nardini et al., 2015; Mizumoto et al., 2019a; Laha et al., 2021; Xrism Collaboration et al., 2025). This setup corresponds to the propagation timescale of a UFO traveling about 1 pc from the central black hole. Observationally, UFOs are typically detected at distances $10^{18}~\text{cm}$ (Tombesi et al., 2012; King and Pounds, 2015; Zubovas and King, 2019), placing the 1 pc scale within the observed distribution range.
The wind and ISM parameters are summarized in Tab. 1. The mass outflow rate is set to $\dot{M}\sim 0.01$$1~M_{\odot}/\mathrm{yr}$ (Tombesi et al., 2012; Gofford et al., 2015; Mizumoto et al., 2019b; Laha et al., 2021).
II.2.3 Wind Region